# GLL: A Differentiable Graph Learning Layer for Neural Networks

### v1

@@ 2.3 Classical Graph-based Learning
Let ${\mathcal{X}}=\{x_{1},x_{2},\dots,x_{n}\}\in\mathbb{R}^{d}$ be a set of $d$-dimensional feature vectors. Let $\mathcal{L}\subset{\mathcal{X}}$ be the set of labeled indices that identifies which feature vectors have labels $y_{i}\in\{1,2,\dots,k\}$ (for a $k$-class classification problem). The task of semi-supervised learning is to infer the labels on the unlabeled index set $\mathcal{U}={\mathcal{X}}\setminus\mathcal{L}$. The labels $y_{i}$ are represented by one-hot vectors; that is, the vector $e_{y_{i}}\in\mathbb{R}^{k}$ of all zeros, except a 1 in the $y_{i}^{th}$ entry. We use the geometric structure of a similarity graph constructed on ${\mathcal{X}}$ to make predictions. Using a symmetric similarity function $w:{\mathcal{X}}\times{\mathcal{X}}\to\mathbb{R}^{+}$, we construct the similarity graph’s weight matrix $W$, where $w_{x_{i}x_{j}}$ or $w_{ij}$ is the edge weight between $x_{i}$ and $x_{j}$ ($w_{xy}$ denotes generic elements $x,y\in{\mathcal{X}}$). The weight matrix $W$ is symmetric and the edge weights correspond to similarity between nodes, with larger values indicating higher similarity. Because pairwise similarities and dense matrix operations are expensive to compute, $W$ is typically approximated using a sparse or low rank representation. Here we use a $k$-nearest neighbors search on ${\mathcal{X}}$ with $k\ll n$ so that $W$ is sparse (see Section 2.1).
Laplace learning (Zhu et al. (2003)), also called label propagation, is a seminal method in graph-based learning. The labels are inferred by solving the graph Laplace equation
$\Delta u(x)&=0,&&\text{if }x\in\mathcal{U}\\ u(x)&=g(x),&&\text{if }x\in\mathcal{L},\end{aligned}\right.$ (10)
@@
subject to the condition that $u=g$ on the labeled set $\mathcal{L}$. The necessary condition satisfied by any minimizer of (11) is exactly the Laplace equation (10). Furthermore, as long as the graph is connected [^3] In fact, a weaker condition that the graph is connected to the labeled data $\mathcal{L}$ is sufficient, which is equivalent to asking that every connected component of the graph has at least one labeled data point. the solution of (10) exists and is unique, as is the minimizer of (11) (see Calder (2018)).
@@
$\min_{u\in{\ell^{2}({\mathcal{X}})}}\left\{\left\|\nabla u\right\|_{{\ell^{2}({\mathcal{X}}^{2})}}^{2}+\lambda\sum_{x\in\mathcal{L}}\|u(x)-g(x)\|^{2}\right\}.$ (12)
@@
where $\xi(x)=0$ if $x\in\mathcal{U}$ and $\xi(x)=1$ if $x\in\mathcal{L}$. We recover Laplace learning (10) in the limit as $\lambda\to 0$.
@@
There are also a variety of graph-based learning techniques that do not require labeled information, such as the Ginzburg-Landau formulation (Garcia-Cardona et al., 2014)
@@
where $G$ is a many-well potential, with wells located at the one hot label vectors. For binary classification, the wells may be positioned at +1/1+1/-1+ 1 / - 1 and $G(u)=(1-u^{2})^{2}$. The MBO methods pioneered by Merkurjev et al. (2013); Garcia-Cardona et al. (2014) are effective approximations of the Ginzburg-Landau energy (14) when $\lambda\gg 1$. This method was first developed in a semi-supervised setting but also used for modularity optimization Boyd et al. (2018) and a piecewise constant Mumford-Shah model in an unsupervised setting Hu et al. (2015).
@@ 2.4 Diagonal Perturbation
A modification to Laplace learning (Section 2.3) is to add a type of Tikhonov regularization in the energy (see e.g. Miller and Calder (2023)), leading to the minimization problem
@@
subject to the condition that $u=g$ on the labeled set $\mathcal{L}$. Tikhonov regularization can also be added to the soft constrained Laplace learning (12). The minimizer of (15) is the solution of the graph Laplace equation (10) except that the graph Laplacian $\Delta$ is replaced by the diagonal perturbed Laplacian $\Delta_{\tau}=\Delta+\tau I$ (likewise for the soft constrained problem (13)).
This regularization term increases the convergence speed of iterative algorithm by improving the condition number of the graph Laplacian. Moreover, it induces exponential decay away from labeled data points. This inhibits label propagation and weakens predictions further from labeled data in the latent space, which can limit overconfident predictions. In Section 4.1, we show that employing this regularization in GLL causes denser clustering of similar samples in the latent space learned by the neural network, which may be desireable for classification tasks.
@@ 2.5 $p$-Laplace learning
There are also graph-learning methods based on the nonlinear graph $p$-Laplacian (Zhou and Belkin (2011); El Alaoui et al. (2016); Flores et al. (2022); Slepcev and Thorpe (2019); Calder (2018)). One such example is the variational $p$-Laplacian, which stems from solving the optimization problem
@@
subject to $u=g$ on $\mathcal{L}$. Taking values of $p\gg 2$ was proposed in (El Alaoui et al., 2016) as a method for improving Laplace learning in the setting of very few labeled data points, and analysis of the $p$-Laplacian was carried out in Slepcev and Thorpe (2019); Calder (2018). The minimizer of the $p$-variational problem (16) solves the graph $p$-Laplace equation
$\operatorname{div}(\|\nabla u\|^{p-2}\nabla u)(x)&=0,&&\text{if }x\in\mathcal{U}\\ u(x)&=g(x),&&\text{if }x\in\mathcal{L}.\end{aligned}\right.$ (17)
Another form of the $p$-Laplacian—the game theoretic $p$-Laplacian—was proposed and studied in (Calder, 2018; Calder and Drenska, 2024).
@@ 2.6 Poisson learning
Another recently proposed approach for low label rate graph-based semi-supervised learning is Poisson learning (Calder et al., 2020). Instead of solving the graph Laplace equation (10), Poisson learning solves the graph Poisson equation
@@
where $\bar{g}=\frac{1}{m}\sum_{x\in\mathcal{L}}g(x)$, $m$ is the number of labeled examples in $\mathcal{L}$, and $\delta_{x}(y)=0$ if $x\neq y$ and $\delta_{x}(x)=1$. In Poisson learning, the labels are encoded as point sources and sinks in a graph Poisson equation, which improves the performance at very low label rates (Calder et al., 2020). Poisson learning also has the variational interpretation
@@ 3 Automatic differentiation through graph learning
In the backward pass, GLL receives the upstream gradients coming from a loss function and need to return the gradients of the loss with respect to the neural network’s feature representations. The calculations for backpropagation can be divided into two distinct steps: tracking the gradients through the graph Laplace equation with respect to the entries in the weight matrix, which is addressed in Section 3.1, and then tracking the gradients from the weight matrix to the feature vectors, which is addressed in Section 3.2. In practice, we combine the implementation of these two terms within the same autodifferentiation function, which allows for more efficient calculations.
@@ 3.1 Backpropagation through graph Laplace equations
In this section, we present the derivation of the backpropagation equations from the solution of the label propagation problem to the weight matrix. For simplicity, we restrict our attention binary classification, since the multi-class setting of any graph learning algorithm is equivalent to the one-vs-rest approach.
In order to make our results general, we work with a general nonlinear elliptic equation on a graph, which we review now. Given a function $\phi:\mathbb{R}\times{\mathcal{X}}^{2}\to\mathbb{R}$ and a vector field $\mathbf{v}\in{\ell^{2}({\mathcal{X}}^{2})}$, we define the function $\phi(\mathbf{v},\cdot)\in{\ell^{2}({\mathcal{X}}^{2})}$ by
@@
The notion of ellipticity is also called monotonicity in some references (Barles and Souganidis, 1991; Calder and Ettehad, 2022), and has appeared previously in the study of graph partial differential equations (Manfredi et al., 2015; Calder, 2018).
We will consider here a nonlinear graph Laplace-type equation
$\tau(u(x),x)+\operatorname{div}\phi(\nabla u,\cdot)(x)&=f(x),&&\text{if }x\in{\mathcal{X}}\setminus\mathcal{L}\\ u(x)&=g(x),&&\text{if }x\in\mathcal{L},\end{aligned}\right.$ (19)
where $\tau:\mathbb{R}\times{\mathcal{X}}\to\mathbb{R}$, $f\in{\ell^{2}({\mathcal{X}})}$, $\mathcal{L}\subset{\mathcal{X}}$, and $g:{\mathcal{X}}\to\mathbb{R}$. The graph Laplace equation generalizes all of the examples of graph Laplace-based semi-supervised learning algorithms given in Section 2. In particular, we allow for the situation that $\mathcal{L}=\varnothing$. We will assume the equation (19) is uniquely solvable; see Calder and Ettehad (2022) for conditions under which this holds. However, at the moment we do not place any assumptions on $\tau$ and $\phi$; in particular, $\phi$ need not be elliptic, preserve vector fields, or be symmetric.
@@
An example is the graph $p$-Laplace equation for $p\geq 1$, where $\tau(z,x)=\tau z$ for a nonnegative constant $\tau$, $f\in{\ell^{2}({\mathcal{X}})}$, and
@@
As discussed previously, we have in mind that the graph PDE (19) is a block within a neural network that is trained end to end. The inputs to the graph learning block are any one of the following objects (or all of them):
  • The weight matrix entries $w_{xy}$.
  • The source term $f$.
  • The boundary condition $g$.
In all cases, the output of the graph learning block is the solution $u\in{\ell^{2}({\mathcal{X}})}$. Let us keep in mind that, while we have used functional analysis notation to illuminate the main ideas more clearly, the inputs and outputs of the block are all vectors in Euclidean space; i.e., $u\in{\ell^{2}({\mathcal{X}})}$ can be identified as a vector in $\mathbb{R}^{n}$ where $n$ is the number of nodes in the graph.
$\nabla_{u}\mathcal{J}(x)=\partial_{u(x)}\mathcal{J},$
$\nabla_{g}\mathcal{J}(z)$ $=\left\langle\partial_{g(x)}u,\nabla_{u}\mathcal{J}\right\rangle_{{\ell^{2}({\mathcal{X}})}},$ (22)
for all $x,y\in{\mathcal{X}}$ and $z\in\mathcal{J}$. The quantities $\nabla_{W}\mathcal{J}$, $\nabla_{f}\mathcal{J}$, and $\nabla_{g}\mathcal{J}$ are the outputs of the graph learning autodifferentiation procedure, and are then fed into the preceding autodifferentiation block. Letting $m$ be the number of labeled points, $\nabla_{W}\mathcal{J}$ is an $n\times n$ matrix, while $\nabla_{f}\mathcal{J}$ is a length $n-m$ vector and $\nabla_{g}\mathcal{J}$ is a length $m$ vector.
@@
$\tau_{z}(u(x),x)v(x)+\operatorname{div}\left(\phi_{q}(\nabla u,\cdot)\nabla v\right)&=\nabla_{u}\mathcal{J}(x),&&\text{if }x\in{\mathcal{X}}\setminus\mathcal{L}\\ v(x)&=0,&&\text{if }x\in\mathcal{L}\end{aligned}\right.$ (23)
for the unknown function $v\in{\ell^{2}({\mathcal{X}})}$. Note that the adjoint equation (23) depends on the solution $u$ of (19), which was computed in the forward pass of the neural network and can be saved for use in backpropagation. We also note that, even though the graph PDE (19) is nonlinear, the adjoint equation (23) is always a linear equation for $v$.
@@
Assume $z\mapsto\tau(z,x)$ and $q\mapsto\phi(q,x,y)$ are continuously differentiable for all $x,y\in{\mathcal{X}}$. Let $f\in{\ell^{2}({\mathcal{X}})}$, $g:\mathcal{L}\to\mathbb{R}$, and suppose $u\in{\ell^{2}({\mathcal{X}})}$ satisfies (19). If the adjoint equation (23) admits a unique solution $v\in{\ell^{2}({\mathcal{X}})}$, then the partial derivatives $\partial_{w_{xy}}u,\partial_{f(x)}u,\partial_{g(x)}u\in{\ell^{2}({\mathcal{X}})}$ exist and are continuous in the variables $w_{xy},f(x)$ and $g(x)$, respectively. Furthermore, it holds that
@@
$\nabla_{g}\mathcal{J}(x)$ $=\nabla_{u}\mathcal{J}(x)-\tau_{z}(u(x),x)v(x)-\operatorname{div}\left(\phi_{q}(\nabla u,\cdot)\nabla v\right)(x)\ \ \text{for}\ x\in\mathcal{L}.$ (26)
@@
Now, the adjoint equation is utilized throughout the whole proof in a similar way. Given $h\in{\ell^{2}({\mathcal{X}})}$ with $h(x)=0$ for $x\in\mathcal{L}$, we can compute the inner produce with $\nabla_{u}\mathcal{J}$ as follows:
$\left\langle h,\nabla_{u}\mathcal{J}\right\rangle_{{\ell^{2}({\mathcal{X}})}}=\left\langle\tau_{z}(u,\cdot)h+\operatorname{div}\left(\phi_{q}(\nabla u,\cdot)\nabla h\right),v\right\rangle_{{\ell^{2}({\mathcal{X}})}}.$ (27)
We will use this identity multiple times below. To see that (27) holds, we compute
$\left\langle h,\nabla_{u}\mathcal{J}\right\rangle_{{\ell^{2}({\mathcal{X}})}}$ $=\left\langle h,\tau_{z}(u,\cdot)v+\operatorname{div}\left(\phi_{q}(\nabla u,\cdot)\nabla v\right)\right\rangle_{{\ell^{2}({\mathcal{X}})}}$
$=\left\langle h,\tau_{z}(u,\cdot)v\right\rangle_{{\ell^{2}({\mathcal{X}})}}+\left\langle h,\operatorname{div}\left(\phi_{q}(\nabla u,\cdot)\nabla v\right)\right\rangle_{{\ell^{2}({\mathcal{X}})}}$
$=\left\langle\tau_{z}(u,\cdot)h,v\right\rangle_{{\ell^{2}({\mathcal{X}})}}+\left\langle\nabla h,\phi_{q}(\nabla u,\cdot)\nabla v\right\rangle_{{\ell^{2}({\mathcal{X}}^{2})}}$
$=\left\langle\tau_{z}(u,\cdot)h,v\right\rangle_{{\ell^{2}({\mathcal{X}})}}+\left\langle\phi_{q}(\nabla u,\cdot)\nabla h,\nabla v\right\rangle_{{\ell^{2}({\mathcal{X}}^{2})}}$
$=\left\langle\tau_{z}(u,\cdot)h,v\right\rangle_{{\ell^{2}({\mathcal{X}})}}+\left\langle\operatorname{div}\left(\phi_{q}(\nabla u,\cdot)\nabla h\right),v\right\rangle_{{\ell^{2}({\mathcal{X}})}}$
$=\left\langle\tau_{z}(u,\cdot)h+\operatorname{div}\left(\phi_{q}(\nabla u,\cdot)\nabla h\right),v\right\rangle_{{\ell^{2}({\mathcal{X}})}}.$
@@
$\tau_{z}(u(x),x)h(x)+\operatorname{div}\left(\phi_{q}(\nabla u,\cdot)\nabla h\right)=-\frac{1}{2}\delta_{r}(x)(\phi(\nabla u(r,s),r,s)-\phi(\nabla u(s,r),s,r))$ (28)

### v2

@@ 2.3 Classical Graph-based Learning
Let ${\mathcal{X}}=\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{n}\}\in\mathbb{R}^{d}$ be a set of $d$-dimensional feature vectors, and let $W$ denote a corresponding weight matrix, which may be constructed as a similarity matrix, as discussed in Section 2.1. This endows the data set with a graph structure. Let $\mathcal{L}\subset{\mathcal{X}}$ be the set of labeled indices that identifies which feature vectors have labels $y_{i}\in\{1,2,\dots,k\}$ (for a $k$-class classification problem). The task of graph-based semi-supervised learning is to utilize the graph structure to infer the labels on the unlabeled index set $\mathcal{U}={\mathcal{X}}\setminus\mathcal{L}$. The labels $y_{i}$ are represented by one-hot vectors; that is, the vector $e_{y_{i}}\in\mathbb{R}^{k}$ of all zeros, except a 1 in the $y_{i}^{th}$ entry. As in Section 2.2, generic elements of $\mathcal{X}$}, i.e., graph vertices, are denoted as $x,y\in{\mathcal{X}}$, and the corresponding entry of the weight matrix $W$ is denoted $w_{xy}$.
Laplace learning (zhu2003semi), also called label propagation, is a seminal method in graph-based learning. The labels are inferred by solving the graph Laplace equation
$\left\{\begin{aligned} \Delta u(x)&=0,&&\text{if }x\in\mathcal{U}\\ u(x)&=g(x),&&\text{if }x\in\mathcal{L},\end{aligned}\right$. (10)
@@
subject to the condition that $u=g$ on the labeled set $\mathcal{L}$. The necessary condition satisfied by any minimizer of (11) is exactly the Laplace equation (10). Furthermore, as long as the graph is connected [^3] In fact, a weaker condition that the graph is connected to the labeled data $\mathcal{L}$ is sufficient, which is equivalent to asking that every connected component of the graph has at least one labeled data point. the solution of (10) exists and is unique, as is the minimizer of (11) (see calder2018game).
@@
$\min_{u\in{\ell^{2}({\mathcal{X}})}}\left\{\left\|\nabla u\right\|_{{\ell^{2}({\mathcal{X}}^{2})}}^{2}+\lambda\sum_{x\in\mathcal{L}}\|u(x)-g(x)\|^{2}\right\},$ (12)
@@
where $\xi(x)=0$ if $x\in\mathcal{U}$ and $\xi(x)=1$ if $x\in\mathcal{L}$. We recover Laplace learning (10) in the limit as $\lambda\to\infty$.
@@
There are also a variety of graph-based learning techniques that do not require labeled information, such as the Ginzburg-Landau formulation (garcia2014multiclass)
@@
where $G$ is a many-well potential, with wells located at the one hot label vectors. For binary classification, the wells may be positioned at $\pm 1$ and $G(u)=(1-u^{2})^{2}$. The MBO methods pioneered by merkurjev_mbo_2013; garcia2014multiclass are effective approximations of the Ginzburg-Landau energy (14) when $\lambda\gg 1$. This method was first developed in a semi-supervised setting but also used for modularity optimization Boyd18 and a piecewise constant Mumford-Shah model in an unsupervised setting Mumfordgraph.
@@ 2.4 Diagonal Perturbation
A modification to Laplace learning (Section 2.3) is to add a type of Tikhonov regularization in the energy (see e.g. miller2023poisson), leading to the minimization problem
@@
subject to the condition that $u=g$ on the labeled set $\mathcal{L}$. Tikhonov regularization can also be added to the soft constrained Laplace learning (12). The minimizer of (15) is the solution of the graph Laplace equation (10) except that the graph Laplacian, $\Delta$, is replaced by the diagonal perturbed Laplacian $\Delta_{\tau}=\Delta+\tau I$ (likewise for the soft constrained problem (13)).
This regularization term increases the convergence speed of iterative algorithm by improving the condition number of the graph Laplacian. Moreover, it induces exponential decay away from labeled data points. This inhibits label propagation and weakens predictions further from labeled data in the latent space, which can limit overconfident predictions. In Section 4.2, we show that employing this regularization in the GLL causes denser clustering of similar samples in the latent space learned by the neural network, which may be desireable for classification or clustering tasks.
@@ 2.5 $p$-Laplace learning
There are also graph-learning methods based on the nonlinear graph $p$-Laplacian (zhou2011semi; el2016asymptotic; flores2022analysis; slepcev2019analysis; calder2018game). One such example is the variational $p$-Laplacian, which stems from solving the optimization problem
@@
subject to $u=g$ on $\mathcal{L}$. Taking values of $p\gg 2$ was proposed in (el2016asymptotic) as a method for improving Laplace learning in the setting of very few labeled data points, and analysis of the $p$-Laplacian was carried out in slepcev2019analysis; calder2018game. The minimizer of the $p$-variational problem (16) solves the graph $p$-Laplace equation
$\left\{\begin{aligned} \operatorname{div}(\|\nabla u\|^{p-2}\nabla u)(x)&=0,&&\text{if }x\in\mathcal{U}\\ u(x)&=g(x),&&\text{if }x\in\mathcal{L}.\end{aligned}\right$. (17)
Another form of the $p$-Laplacian—the game theoretic $p$-Laplacian—was proposed and studied in (calder2018game; calder2024consistency).
@@ 2.6 Poisson learning
Another recently proposed approach for low label rate graph-based semi-supervised learning is Poisson learning (calder2020poisson). Instead of solving the graph Laplace equation (10), Poisson learning solves the graph Poisson equation
@@
where $\bar{g}=\frac{1}{m}\sum_{x\in\mathcal{L}}g(x)$, $m$ is the number of labeled examples in $\mathcal{L}$, and $\delta_{x}(y)=0$ if $x\neq y$ and $\delta_{x}(x)=1$. In Poisson learning, the labels are encoded as point sources and sinks in a graph Poisson equation, which improves the performance at very low label rates (calder2020poisson). Poisson learning also has the variational interpretation
@@ 3 Automatic differentiation through graph learning
In the backward pass, the GLL receives the upstream gradients coming from a loss function and need to return the gradients of the loss with respect to the neural network’s feature representations. The calculations for backpropagation can be divided into two distinct steps: tracking the gradients through the graph Laplace equation with respect to the entries in the weight matrix, which is addressed in Section 3.1, and then tracking the gradients from the weight matrix to the feature vectors, which is addressed in Section 3.2. In practice, we combine the implementation of these two terms within the same autodifferentiation function, which allows for more efficient calculations.
@@ 3.1 Backpropagation through graph Laplace equations
In this section, we present the derivation of the backpropagation equations for computing the gradient of the solution of a graph Laplace equation for label propagation with respect to the ingredients of the equation, including the weight matrix used to construct the graph, and any source terms or boundary conditions that are present in the equation. For simplicity, we restrict our attention to binary classification, since the multi-class setting of any graph learning algorithm is equivalent to the one-vs-rest approach.
3.1.1 Linear graph Laplace equations
In order to elucidate our approach in a simpler setting, we will first present our main results for the linear graph Laplace equation
$\left\{\begin{aligned} \tau u(x)+\Delta u(x)&=f(x),&&\text{if }x\in\mathcal{U}\\ u(x)&=g(x),&&\text{if }x\in\mathcal{L}.\end{aligned}\right$. (19)
Here, $\tau\geq 0$ is a parameter, $\Delta$ is the graph Laplacian, defined in (9), $g$ is the boundary condition, and $f$ is the source term. Depending on the context, the label information can be encoded into the source term $f$ or the boundary condition $g$.
As discussed previously, we have in mind that a graph PDE like (19) is a layer within a neural network that is trained end to end. The inputs to the graph learning layer are any one of the following objects (or all of them):
  • The weight matrix entries $w_{xy}$.
  • The source term $f$.
  • The boundary condition $g$.
In all cases, the output of the graph learning layer is the solution $u\in{\ell^{2}({\mathcal{X}})}$, which feeds into the next layer of the network. Let us keep in mind that, while we have used functional analysis notation to illuminate the main ideas more clearly, the inputs and outputs of the block are all vectors in Euclidean space; i.e., a function $u\in{\ell^{2}({\mathcal{X}})}$ is simply a vector in $\mathbb{R}^{n}$ where $n$ is the number of nodes in the graph.
$\nabla_{u}\mathcal{J}(x)=\partial_{u(x)}\mathcal{J},$
$\nabla_{g}\mathcal{J}(x)$ $=\left\langle\partial_{g(x)}u,\nabla_{u}\mathcal{J}\right\rangle_{{\ell^{2}({\mathcal{X}})}},$ (22)
for all $x,y\in{\mathcal{X}}$. The quantities $\nabla_{W}\mathcal{J}$, $\nabla_{f}\mathcal{J}$, and $\nabla_{g}\mathcal{J}$ are the outputs of the graph learning autodifferentiation procedure and are then fed into the preceding autodifferentiation block. Letting $m$ be the number of labeled points, $\nabla_{W}\mathcal{J}$ is an $n\times n$ matrix, while $\nabla_{f}\mathcal{J}$ is a length nmn-m vector and $\nabla_{g}\mathcal{J}$ is a length $m$ vector.
In the special case of the linear graph Laplace equation (19), backpropgation through the graph learning layer requires solving the adjoint equation
$\left\{\begin{aligned} \tau v(x)+\Delta v(x)&=\nabla_{u}\mathcal{J},&&\text{if }x\in\mathcal{U}\\ v(x)&=0,&&\text{if }x\in\mathcal{L}\end{aligned}\right$. (23)
for the unknown function $v\in{\ell^{2}({\mathcal{X}})}$. Since the left hand side of (19) is self-adjoint, the adjoint equation (23) is nearly identical to the original equation (19) — indeed, the main difference is that the boundary conditions have been set to zero, and the source term is replaced by the upstream gradient $\nabla_{u}\mathcal{J}$, which is the input to the automatic differentiation method. Hence, the computational complexity of the adjoint equation is similar to the original graph Laplace equation that was solved in the forward pass.
In this case, the adjoint equation (23) is uniquely solvable for $v$, provided the graph is connected, and the downstream gradients we wish to compute can all be expressed in terms of $v$ as follows (see Theorem 5):
$\nabla_{W}\mathcal{J}(x,y)$ $=-\nabla u(x,y)v(x)=-(u(x)-u(y))v(x)\ \ \text{for}\ x,y\in{\mathcal{X}},$ (24)
$\nabla_{g}\mathcal{J}(x)$ $=\nabla_{u}\mathcal{J}(x)-\Delta v(x)\ \ \text{for}\ x\in\mathcal{L}.$ (26)
We note that the last expression for $\nabla_{g}\mathcal{J}(x)$ cannot be simplified using the adjoint equation, since it is evaluated on the labeled nodes $\mathcal{L}$ where the adjoint equation does not hold. We also note that it may seem a little strange at first that $\nabla_{W}\mathcal{J}(x,y)$ is not symmetric in $x$ and $y$. However, as we show in the next section, since the weight matrix $W$ is symmetric, we can replace $\nabla_{W}\mathcal{J}(x,y)$ by the symmetrized expression
$\nabla u(x,y)\nabla v(x,y)=-(u(x)-u(y))(v(x)-v(y))$
without affecting any downstream gradient computations. We use this symmetrized expression in practice to simplify computations.
The proofs of the claims in this section are all provided in a more general setting in Theorem 5 in the following section.
3.1.2 Nonlinear elliptic equations on graphs
In order to make our results general - applicalbe to any of the graph-based learning algorithms discussed in Section 2, including nonlinear graph-Laplace equations - we proceed in this section with generalizing the simple graph Laplace equation (19) to a general nonlinear elliptic equation on a graph. Given a function $\phi:\mathbb{R}\times{\mathcal{X}}^{2}\to\mathbb{R}$ and a vector field $\mathbf{v}\in{\ell^{2}({\mathcal{X}}^{2})}$, we define the function $\phi(\mathbf{v},\cdot)\in{\ell^{2}({\mathcal{X}}^{2})}$ by
@@ 3.1.2 Nonlinear elliptic equations on graphs
The notion of ellipticity is also called monotonicity in some references (barles1991convergence; calder2022hamilton), and has appeared previously in the study of graph partial differential equations (manfredi2015nonlinear; calder2018game).
We will consider here a nonlinear graph Laplace-type equation of the form
$\left\{\begin{aligned} \sigma(u(x),x)+\operatorname{div}\phi(\nabla u,\cdot)(x)&=f(x),&&\text{if }x\in{\mathcal{X}}\setminus\mathcal{L}\\ u(x)&=g(x),&&\text{if }x\in\mathcal{L},\end{aligned}\right$. (27)
where $\sigma:\mathbb{R}\times{\mathcal{X}}\to\mathbb{R}$, $f\in{\ell^{2}({\mathcal{X}})}$, $\mathcal{L}\subset{\mathcal{X}}$, and $g:{\mathcal{X}}\to\mathbb{R}$. The graph Laplace equation generalizes all of the examples of graph Laplace-based semi-supervised learning algorithms given in Section 2. In particular, we allow for the situation that $\mathcal{L}=\varnothing$. We will assume the equation (27) is uniquely solvable; see calder2022hamilton for conditions under which this holds. However, at the moment we do not place any assumptions on $\sigma$ and $\phi$; in particular, $\phi$ need not be elliptic, preserve vector fields, or be symmetric.
@@
An example is the graph $p$-Laplace equation for $p\geq 1$, where $\sigma(z,x)=\tau z$ for a nonnegative constant $\tau$, $f\in{\ell^{2}({\mathcal{X}})}$, and
@@
$\left\{\begin{aligned} \sigma_{z}(u(x),x)v(x)+\operatorname{div}\left(\phi_{q}(\nabla u,\cdot)\nabla v\right)&=\nabla_{u}\mathcal{J}(x),&&\text{if }x\in{\mathcal{X}}\setminus\mathcal{L}\\ v(x)&=0,&&\text{if }x\in\mathcal{L}\end{aligned}\right$. (28)
for the unknown function $v\in{\ell^{2}({\mathcal{X}})}$. Recall that $\sigma_{z}$ denotes the partial derivative of the function $(z,x)\mapsto\sigma(z,x)$ in the first variable $z$, i.e., if $\sigma(z,x)=z^{2}$ then $\sigma_{z}(z,x)=2z$. We also note that using the definition of divergence given in (8), the divergence term $\operatorname{div}\left(\phi_{q}(\nabla u,\cdot)\nabla v\right)$ evaluated at a vertex $x\in{\mathcal{X}}\setminus\mathcal{L}$ in (28) can be expressed as
$\frac{1}{2}\sum_{y\in{\mathcal{X}}}w_{xy}(\phi_{q}(\nabla u(x,y),x,y)(v(x)-v(y))-\phi_{q}(\nabla u(y,x),y,x)(v(y)-v(x))).$ (29)
If $\phi$ is symmetric, as per Definition 3, then $\phi_{q}(-q,y,x)=\phi_{q}(q,x,y)$ and this expression can be simplified to read
$\operatorname{div}\left(\phi_{q}(\nabla u,\cdot)\nabla v\right)=\sum_{y\in{\mathcal{X}}}w_{xy}\,\phi_{q}(\nabla u(x,y),x,y)(v(x)-v(y)).$ (30)
This expression is a graph Laplacian applied to the function $v$, with weights that depend on the gradient of $u$. In fact, the first term $\sigma_{z}(u(x),x)v(x)$ in the adjoint equation (28) also depends on the solution $u$ of (27), which was computed in the forward pass of the neural network and can be saved for use in backpropagation. We also note that, even though the graph PDE (27) is nonlinear, the adjoint equation (28) is always a linear equation for $v$.
@@
Assume $z\mapsto\sigma(z,x)$ and $q\mapsto\phi(q,x,y)$ are continuously differentiable for all $x,y\in{\mathcal{X}}$. Let $f\in{\ell^{2}({\mathcal{X}})}$, $g:\mathcal{L}\to\mathbb{R}$, and suppose $u\in{\ell^{2}({\mathcal{X}})}$ satisfies (27). If the adjoint equation (28) admits a unique solution $v\in{\ell^{2}({\mathcal{X}})}$, then the partial derivatives $\partial_{w_{xy}}u,\partial_{f(x)}u,\partial_{g(x)}u\in{\ell^{2}({\mathcal{X}})}$ exist and are continuous in the variables $w_{xy},f(x)$ and $g(x)$, respectively. Furthermore, it holds that
@@
$\nabla_{g}\mathcal{J}(x)$ $=\nabla_{u}\mathcal{J}(x)-\operatorname{div}\left(\phi_{q}(\nabla u,\cdot)\nabla v\right)(x)\ \ \text{for}\ x\in\mathcal{L}.$ (33)
@@
Now, the adjoint equation is utilized throughout the whole proof in a similar way. Given $h\in{\ell^{2}({\mathcal{X}})}$ with $h(x)=0$ for $x\in\mathcal{L}$, we can compute the inner product with $\nabla_{u}\mathcal{J}$ as follows:
$\left\langle h,\nabla_{u}\mathcal{J}\right\rangle_{{\ell^{2}({\mathcal{X}})}}=\left\langle\sigma_{z}(u,\cdot)h+\operatorname{div}\left(\phi_{q}(\nabla u,\cdot)\nabla h\right),v\right\rangle_{{\ell^{2}({\mathcal{X}})}}.$ (34)
To see this we compute
$\left\langle h,\nabla_{u}\mathcal{J}\right\rangle_{{\ell^{2}({\mathcal{X}})}}$ $=\left\langle h,\sigma_{z}(u,\cdot)v+\operatorname{div}\left(\phi_{q}(\nabla u,\cdot)\nabla v\right)\right\rangle_{{\ell^{2}({\mathcal{X}})}}$
$=\left\langle h,\sigma_{z}(u,\cdot)v\right\rangle_{{\ell^{2}({\mathcal{X}})}}+\left\langle h,\operatorname{div}\left(\phi_{q}(\nabla u,\cdot)\nabla v\right)\right\rangle_{{\ell^{2}({\mathcal{X}})}}$
$=\left\langle\sigma_{z}(u,\cdot)h,v\right\rangle_{{\ell^{2}({\mathcal{X}})}}+\left\langle\nabla h,\phi_{q}(\nabla u,\cdot)\nabla v\right\rangle_{{\ell^{2}({\mathcal{X}}^{2})}}$
$=\left\langle\sigma_{z}(u,\cdot)h,v\right\rangle_{{\ell^{2}({\mathcal{X}})}}+\left\langle\phi_{q}(\nabla u,\cdot)\nabla h,\nabla v\right\rangle_{{\ell^{2}({\mathcal{X}}^{2})}}$
$=\left\langle\sigma_{z}(u,\cdot)h,v\right\rangle_{{\ell^{2}({\mathcal{X}})}}+\left\langle\operatorname{div}\left(\phi_{q}(\nabla u,\cdot)\nabla h\right),v\right\rangle_{{\ell^{2}({\mathcal{X}})}}$
$=\left\langle\sigma_{z}(u,\cdot)h+\operatorname{div}\left(\phi_{q}(\nabla u,\cdot)\nabla h\right),v\right\rangle_{{\ell^{2}({\mathcal{X}})}}.$
We will use the identity (34) multiple times in what follows.
@@
$\sigma_{z}(u(x),x)h(x)+\operatorname{div}\left(\phi_{q}(\nabla u,\cdot)\nabla h\right)=-\frac{1}{2}\delta_{r}(x)(\phi(\nabla u(r,s),r,s)-\phi(\nabla u(s,r),s,r))$ (35)